OpenStudy (anonymous):

Can someone please help me solve this Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2))

OpenStudy (blank ):

It is -7. You need the steps?

OpenStudy (anonymous):

wat

OpenStudy (blank ):

f(x)=x+8 g(x)=x^(2)-6x-7 f(g(2)) Evaluate f(g(x)) by substituting in the value of g into f. f(x^(2)-6x-7)=x^(2)-6x-7+8 Add 8 to -7 to get 1. f(x^(2)-6x-7)=x^(2)-6x+1 Evaluate the result function by replacing x with 2 in the function. (2)^(2)-6(2)+1 Expand the exponent (2) to the expression. (2^(2))-6(2)+1 Squaring a number is the same as multiplying the number by itself (2*2). (2*2)-6(2)+1 Multiply 2 by 2 to get 4. (4)-6(2)+1 Multiply -6 by the 2 inside the parentheses. (4)-6*2+1 Multiply -6 by 2 to get -12. 4-12+1 Subtract 12 from 4 to get -8. -8+1 Add 1 to -8 to get -7. -7 Hope that helps.